What is the most a rational investor should be willing to pay for the right to receive $12,000 per year (the first payment is at the end of the first year) for 20 years if her annual opportunity cost is 12%?  (Choose the closest answer)
Answer #1

This question requires application of PV of annuity formula, according to which

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

For this question,

PV = 12000* 1- (1 +0.12) -20 PV = 12000* 0.12

PV = 20000 * 7.469444

PV = $89,633.32